A 1.00 kg sample of ethylene glycol (C2H6O2) is added to 4450 g of water in a car radiator What is the boiling point of the resulting solution? Kb of water is 0.512 OC kg mori, Tb for water 100.0 oC.
Answer #1

∆Tb = 1000×Kb×w/W×m

W is mass of solvent, w is mass of solute ,m is molecular mass of solute.

∆Tb =(Tf-100) =1000×0.512×1/4.450×62

Tf-100 = 1.98

Tf = 100+1.98 =101.98℃